Evaluate each series to the given term.
2+4+6+8+...10th term

Answer by mananth(16946)

it is an Arithmetic sequence
a=2, d=2
t10 = a+(n-1)d
t10 = 2+(10-1)2
t10 =2+18
=20
10th term =20